OpenStudy (anonymous):

Suppose you know there are originally 2 red marbles in the bag. How many marbles are there in all? Show your work.

OpenStudy (smurfy14):

not enough info

OpenStudy (anonymous):

The probability of randomly selecting a red marble from a bag, replacing it, and selecting a red marble again is 1/25

OpenStudy (gw2011):

The probability of getting a red marble is 2/x where x is the total number of marbles. Then the probability of randomly selecting a red marble, replacing it, and selecting a red marble again is (2/x)(2/x). Therefore, (2/x)(2/x)=1/25 4/x^2=1/25 4=(1/25)(x^2) (4)(25)=x^2 100=x^2 10=x Therefore, there are 10 marbles in all